OpenStudy (anonymous):

PLEASE HELP!!!!!!!!! What is the center and radius of the following circle: x2 + (y – 6)2 = 50 @jigglypuff314

jigglypuff314 (jigglypuff314):

circle formula (x−h)^2+(y−k)^2=r^2 when (h, k) is center and r= radius

OpenStudy (anonymous):

look the general eq of any circle is \[(x-a)^{2}+(y-b)^{2}=r ^{2}\]

OpenStudy (anonymous):

where (a,b)=the center coord. r=radius

OpenStudy (anonymous):

JIggly I know that...But I just need help with finding "h"

jigglypuff314 (jigglypuff314):

so x2 + (y – 6)2 = 50 you can look at as (x-0)^2 + (y-6)^2 = (root50)^2 so h = 0 k = 6 and r = root50

OpenStudy (anonymous):

r=7.02 as well? Would that work?

jigglypuff314 (jigglypuff314):

7.07 if you want...

OpenStudy (anonymous):

That's great! Thanks :)